OpenStudy (anonymous):

A city has a population of 370,000 people. Suppose that each year the population grows by 6.25%. What will the population be after 10 years?

OpenStudy (anonymous):

OKAY! So i thought I got the hand of this but I DON'T :(

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

same as the last one

OpenStudy (misty1212):

what is \(100\%+6.25\%\)?

OpenStudy (anonymous):

The first problem we should look at deals with population growth. If we look at what makes a population grow over one year maybe we can develop a model that will help us predict the future. Suppose that the population of a certain town is P is increasing at a constant rate of R each year. We can see that the population of our town one year later will be: P + P*R = P*( 1 + R ) Two years later the population of our town will be: P*( 1 + R ) + (P*( 1 + R ))*R = P ( 1 + R ) ^ 2 So, in general the model of population growth looks like: P*( 1 + R )^ (t-1) + (P*( 1 + R )^ (t-1))*R) = P*( 1 + R ) ^ t , where t is the number of years later.

OpenStudy (misty1212):

??`

OpenStudy (anonymous):

106.25

OpenStudy (misty1212):

ok good

OpenStudy (misty1212):

so do \[370,000\times (1.0625)^{10}\]

OpenStudy (anonymous):

678408? I think I know what I've been doing wrong now..

OpenStudy (misty1212):

probably put it in the calculator wrong is my guess

OpenStudy (misty1212):

i would use this http://www.wolframalpha.com/input/?i=370000%281.0625%29%5E10

OpenStudy (anonymous):

How would I solve it if it was asking about decreasing rather than growing? @misty1212

OpenStudy (misty1212):

decreasing by say \(2\%\) the compute \(100\%-2\%=98\%=.98\)

OpenStudy (anonymous):

NOW I GET IT! THANKS <3